OpenStudy (anonymous):

What is the slope of the line passing through the points (–3, 4) and (2, –1)?

OpenStudy (nurali):

Same

OpenStudy (anonymous):

c?

OpenStudy (nurali):

No u solve and give me answer.

OpenStudy (nurali):

Formula m=y2-y1/x2-x1

OpenStudy (anonymous):

b

OpenStudy (nurali):

what b?

OpenStudy (anonymous):

a

OpenStudy (nurali):

a?

OpenStudy (anonymous):

d

OpenStudy (nurali):

where is option

OpenStudy (anonymous):

ohhh srry

OpenStudy (anonymous):

a)-8/5 b)-5/8 c)5/8 d)8/5

OpenStudy (nurali):

Wrong choice.

OpenStudy (anonymous):

i think its b

OpenStudy (nurali):

All option are wrong.

OpenStudy (anonymous):

omg yes i put down wrong answers soooo srry

OpenStudy (nurali):

No Problem

OpenStudy (anonymous):

a)1 b)-5/3 c)3/5 d)-1

OpenStudy (nurali):

Ok now tell me the correct one.

OpenStudy (anonymous):

c

OpenStudy (nurali):

wrong

OpenStudy (nurali):

u try to solve

OpenStudy (nurali):

m=y2-y1/x2-x1

OpenStudy (nurali):

(x1,y1)=(-3,4) (x2,y2)=(2,-1)

OpenStudy (nurali):

m=-1-4/2+5